### v1

# Fate of transient order parameter domain walls in ultrafast experiments

Abstract
In ultrafast experiments, an optical pump pulse often generates transient domain walls of the order parameter in materials with spontaneous symmetry breaking, due to either a finite penetration depth of the light on a three-dimensional (3D) material, or a finite spot size on a two-dimensional (2D) material. We clarify the decaying process of such a domain wall that is caused by fluctuations of the order parameters. We study a generic system with $U(1)$-symmetric order, and those with an additional weak $Z_{2}$ ($U(1)$-symmetry-breaking) term, representing the charge-density-wave (CDW) orders in recent experiments. The decay process comprises two non-trivial stages. During the first stage, exponentially growing thermal fluctuations convert the domain wall into an interface with randomly distributed topological defects. In the second stage, the topological defects undergo a coarsening dynamics within the interface. For a 2D interface in the 3D system, the coarsening dynamics leads to a diffusive growth of the correlation length. For a one-dimensional (1D) interface in the 2D system with the weak $Z_{2}$ term, the correlation-length growth shows a crossover from diffusive to sub-diffusive behavior. Our theory provides a fundamental physical picture for the dynamics of pump-induced domain walls in ultrafast experiments.
Introduction— Transient manipulation of order parameters by an ultrafast laser pulse has been one of the major topics in condensed matter physics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Recent experiments [25, 26, 27, 28] reported that in several CDW materials, a laser pulse (pump) created domain walls of the CDW order parameter. Thereby, the pump pulse heats the surface of the system, and the CDW order parameter $\psi$ in the surface region is suppressed in time from its initial equilibrium value $\psi_{\text{eq}}$ to zero, following a classical equation of motion. From the principle of inertia, $\psi$ may cross zero, and stabilize in the energy minimum at $-\psi_{\text{eq}}$ after the system is cooled down [25, 26, 27, 28], getting ‘inverted’. Due to the finite penetration depth of the pump pulse, only the order parameter in the surface region gets inverted, yielding a domain wall (white plane in the left panel of Fig. 1) between the surface region and the bulk at time $t=t_{0}$, see Sec. I of Ref. [29]. Within a mean-field theory of the time-dependent Ginzburg-Landau (TDGL) dynamics that neglects fluctuations, the domain wall undergoes no further evolution [25, 26, 28, 27].
In this letter, we propose that due to fluctuations, the domain wall will decay following a two-stage dynamics. In the first stage, the domain wall decays by the exponential growth of unstable fluctuations [30, 16] of $\psi$ on the wall, which happens both in the $U(1)$ symmetric case, and in the system with a weak $Z_{2}$ term. The dynamics is characterized by increasing magnitude and correlation length of the fluctuations. At the end of the first stage, around a timescale $t\simeq t_{\rm c}$, the wall is transformed into an interface containing randomly distributed topological defects, as shown by the vortex strings in Fig. 1 for a 3D system. In 2D systems, the interface is 1D and the defects are vortices and antivortices.
In the second stage, the motion and annihilation of topological defects on the interface dominate the dynamics. It happens under driving forces originating from the tension of the vortex strings in 3D systems and the attraction between vortices and antivortices in 2D systems  [31, 32, 33, 34, 35]. Universal features of the second-stage dynamics are characterized by a mean distance $L(t)$ between the defects on the interface. The system undergoes a self-similar evolution with a global increase of $L(t)$, leading to a coarsening phenomenon [30, 32, 31, 31, 34], as illustrated in Fig. 1. We show that in 3D systems, the coarsening dynamics exhibits a diffusive behavior in both the $U(1)$-symmetric case and the case with the additional weak $Z_{2}$ term, i.e., the length scale increases as $L(t)\propto t^{1/2}$. In 2D systems with the weak $Z_{2}$ term, $L(t)\propto t^{p}$ exhibits a crossover from $p\approx 1/2$ to $p\approx 0$ as $L$ gets larger, namely a diffusive-to-subdiffusive crossover. The length scale $L(t)$ can be measured from the structure factor in time-resolved X-ray diffraction experiments [36, 37].
Figure 1: Illustration of the temporal evolution of a 3D sample. At $t=t_{\text{pump}}$, the surface region is pumped by an ultrafast pulse represented by the red segment on the time axis. An order parameter domain wall is created at $t=t_{0}$ on the white interface. The four planes on the right show the order parameter configuration of the interface at different times. The black curves denote the vortex strings.
$U(1)$-symmetric model— We consider the “model A” dynamics [38, 39], whose mean-field version is also called the ‘TDGL equation’, together with a $U(1)$-invariant free energy:
$\frac{1}{\gamma}\partial_{t}\psi=-\frac{1}{E_{\text{c}}}\frac{\delta F}{\delta\psi^{*}}+\eta\,,$ (1)
$F=E_{\text{c}}\int f\ \ d^{D}\bm{r}\,,\quad f=\left(|\psi|^{2}+\alpha\right)^{2}+\xi_{0}^{2}|\nabla\psi|^{2}\,.$ (2)
Here $\psi(\bm{r},t)=\psi_{1}+i\psi_{2}$ is the dimensionless complex order parameter, $\gamma$ is the relaxation rate, $E_{\text{c}}$ is the condensation energy density, $D$ is the space dimension, $\alpha\sim(T-T_{\text{c}})/T_{\text{c}}$ is negative at a temperature $T$ below the critical temperature $T_{\text{c}}$, and $\xi_{0}$ is the bare coherence length. The thermal noise field $\eta=\eta_{1}+i\eta_{2}$ arises from a thermal bath made of the other degrees of freedom, including high-energy electrons and phonons. We characterize the noise by a correlator $\langle\eta_{i}(\bm{r},t)\eta_{i^{\prime}}(\bm{r^{\prime}},t^{\prime})\rangle=2T(\gamma E_{\text{c}})^{-1}\delta_{ii^{\prime}}\delta(\bm{r}-\bm{r^{\prime}})\delta(t-t^{\prime})$ with an effective temperature $T$ of the bath and $i/i^{\prime}=1,2$. Here the Ginzburg parameter $\zeta=|\alpha|^{D/2-2}T/(E_{\text{c}}\xi_{0}^{D})$ plays the role of a dimensionless measure of the fluctuations in the non-equilibrium process as well as the equilibrium thermal fluctuations  [16, 40]. Eqs. (1)(2) effectively describe the dynamics of the mean-field order parameter and its fluctuations at temperatures close to $T_{\text{c}}$ [38, 40], in incommensurate CDW [14, 15], spin density waves (SDW) [41], excitonic order, and superconductivity [39, 42, 40, 16].
We first consider the 2D domain wall in the 3D systems ($D=3$), where the domain wall in Fig. 1 is set as the $z=0$ plane. An order parameter configuration of the domain wall satisfies the saddle-point condition $\delta F/\delta\psi=0$. Without loss of generality, we choose the real solution $\psi_{0}(\bm{r})=\sqrt{|\alpha|}\tanh(z/\xi)$ as shown by the gray line in Fig. 2(a), where $\xi=\xi_{0}/\sqrt{|\alpha|}$ is the coherence length.
Figure 2: Illustration of the decay of the domain wall. (a) The ‘Mexican hat’ free energy landscape (the potential part) of Eq. (2) plotted on the complex plane of $\psi$. (b,c) The phase of $\psi$ plotted on the $y-z$ plane at (b) $t=t_{0}$ and (c) $t\geq t_{\text{c}}$, showing the domain wall at $z=0$. The $\psi$ and free energy along the gray, black, and white lines are also plotted in (a). In the initial domain-wall configuration, $\psi=\psi_{0}(z)$ changes from $-\sqrt{|\alpha|}$ to $\sqrt{|\alpha|}$ crossing the potential peak along the gray line. To lower the energy, $\psi$ fluctuates in the imaginary direction. After $t_{\text{c}}$, random phase domains appear on the interface. The loop denoted by black and white lines in (c) encloses a vortex with a phase winding $2\pi$. In a 2D interface of a 3D system, the interface becomes full of vortex-strings as shown in Fig. 1.
To study the first stage of dynamics, we set $t_{0}=0$ and write $\psi(\bm{r},t)=\psi_{0}(\bm{r})+\delta\psi(\bm{r},t)$ where $\delta\psi=\phi_{1}+i\phi_{2}$ is the fluctuation field with the real part $\phi_{1}$ and imaginary part $\phi_{2}$. The linearized dynamics of $\delta\psi$ is expanded from Eq. (1) as
$\frac{1}{\gamma}\partial_{t}\phi_{a}$ $=\left(-\hat{L}_{a}+\xi_{0}^{2}\nabla_{\bm{R}}^{2}\right)\phi_{a}+\eta_{a}$ (3)
where $a=1,2$, $\hat{L}_{1}=-\xi_{0}^{2}\partial_{z}^{2}+2\alpha+6\psi_{0}^{2}$, $\hat{L}_{2}=-\xi_{0}^{2}\partial_{z}^{2}+2\alpha+2\psi_{0}^{2}$, and $\nabla_{\bm{R}}^{2}=\partial_{x}^{2}+\partial_{y}^{2}$ is the Laplacian in $D-1$ dimensions. The eigenfunction of $\hat{L}_{a}+\xi_{0}^{2}\nabla_{\bm{R}}^{2}$ can be written as $\phi_{a}^{\bm{K}}(\bm{r})=u_{a}(z)e^{i\bm{K}\cdot\bm{R}}$ with eigenvalue $\epsilon_{a}+\xi_{0}^{2}K^{2}$, where $u_{a}(z)$ is the eigenfunction of $\hat{L}_{a}$ with eigenvalue $\epsilon_{a}$. Here $\bm{R}$($\bm{K}$) denotes the spatial coordinate (momentum) in the $(D-1)$-dimensional plane normal to $z$. Eigenvalues of $\hat{L}_{1}$ are all non-negative, while $\hat{L}_{2}$ has one negative eigenvalue, $\epsilon_{2}=\alpha$, corresponding to a bound-state eigen-function $u_{2}(z)=\text{sech}(z/\xi)$, see Sec. II A of Ref. [29]. It implies that the long-wavelength fluctuations in the imaginary part direction, decomposed as $\phi_{2}^{\bm{K}}(\bm{r})=\text{sech}(z/\xi)e^{i\bm{K}\cdot\bm{R}}$ with $K<\sqrt{|\alpha|}/\xi_{0}$, will grow exponentially in time with the exponent $\left(|\alpha|-\xi_{0}^{2}K^{2}\right)\gamma t$ [16].
With $\phi_{2}(\bm{r})=\varphi(\bm{R},t)\text{sech}(z/\xi)$, the correlation function $C(\bm{R},t)=\langle\varphi(\bm{0},t)\varphi(\bm{R},t)\rangle$ depicts the real space structure of the fluctuation within the interface. From Eq. (3), it has the following analytical expression for $\gamma t\gg 1/|\alpha|$ (see Sec. II B of Ref. [29]):
$C\left(\bm{R},t\right)\approx\zeta|\alpha|\frac{e^{2|\alpha|\gamma t}}{\sqrt{8\pi|\alpha|\gamma t}}I_{0}\left(\frac{R^{2}}{16\xi_{0}^{2}\gamma t}\right)e^{-\frac{R^{2}}{16\xi_{0}^{2}\gamma t}}$ (4)
with the modified Bessel function $I_{0}$ of the first kind. Therefore, the fluctuation magnitude grows exponentially with the correlation length growing as $L(t)=4\xi_{0}\sqrt{\gamma t}$, a diffusive behavior. Note that due to the initial fluctuations $\langle\phi_{2}(\bm{r})\phi_{2}(\bm{r^{\prime}})\rangle\sim 1/|\bm{r}-\bm{r^{\prime}}|$ at long distances, the spatial correlation crossovers from an exponential decay at $R\ll L(t)$ to a power law decay $C\left(\bm{R},t\right)\propto L(t)/R$ at $R\gg L(t)$. The first-stage dynamics ends at a crossover time $t_{\text{c}}$ when the fluctuation becomes so large that the linear approximation in Eq. (3) fails, and the $|\psi|^{4}$ term in Eq. (2) starts to stop its growth. One may estimate this timescale by $\langle|\psi(\bm{r},t_{\text{c}})|^{2}\rangle\sim|\alpha|$, which gives $C(\bm{0},t_{\text{c}})\sim|\alpha|$ and $\gamma t_{\text{c}}\sim(2|\alpha|)^{-1}\ln(\zeta^{-1})$. Note that $\gamma t_{\text{c}}$ thus obtained is much larger than $1/|\alpha|$ given that $\zeta\ll 1$.
At $t\simeq t_{\text{c}}$, the interface consists of randomly distributed positive-$\phi_{2}$ and negative-$\phi_{2}$ domains with the typical domain size of $4\xi_{0}\sqrt{\gamma t_{\text{c}}}$ [16]. Boundaries between the domains quickly relax into stable topological defects with the phase winding ±$\pm 2\pi$, as illustrated in Fig. 2(c), meaning that the interface becomes full of randomly distributed 2D vortex strings as in Fig. 1.
At $t\gtrsim t_{\text{c}}$, the second stage of dynamics takes over, which has been argued to be a coarsening dynamics [30] of the vortex strings. Thereby, distributions of the vortex strings at different times show a statistical self-similarity, which is characterized by a growing correlation length $L(t)$, or the typical inter-string distance. The coarsening dynamics with $L(t)\propto t^{p}$ and $p>1/2$, $p=1/2$, and $p<1/2$ is referred to as superdiffusive, diffusive, and subdiffusive dynamics, respectively. Eq. (1) implies the equation of motion for the length scale:
$\lambda(L)\frac{dL}{dt}=F_{\text{d}}(L)\ ,$ (5)
see Refs. [30, 32, 31, 31, 34] and Sec. III B of Ref. [29]. Here $\lambda(L)\frac{dL}{dt}$ and $F_{\text{d}}(L)$ are the frictional force with a friction coefficient $\lambda(L)$ and the driving force acting on a unit length of the vortex string, respectively. Eq. (5) is nothing but a balance equation between the two forces. Consider a circular vortex-string loop with the diameter $L$, a typical topological defect, its energy is $E\sim|\alpha|E_{\text{c}}\xi_{0}^{2}L\ln(L/\xi)$ [30, 29]. Therefore, the driving force that contracts the loop is $F_{\text{d}}(L)\sim L^{-1}dE/dL\sim|\alpha|E_{\text{c}}\xi_{0}^{2}L^{-1}\ln(L/\xi)$ in the large $L$ limit. The friction coefficient $\lambda(L)\sim\gamma^{-1}E_{\text{c}}|\alpha|\ln(L/\xi)$ is derived from the damping term [30, 29], i.e. first-order time derivative term in Eq. (1). Combining these two, we obtain a solution of Eq. (5) in the large-$L$ limit as $L(t)\sim\xi_{0}\sqrt{\gamma t}$, which manifests a diffusive coarsening dynamics.
We numerically simulated Eq. (1) with the domain-wall initial condition and observed the coarsening dynamics. To evaluate the length scale $L(t)$ in the simulation, we count a number $N$ of vortices in the $y$-$z$ planes for each discrete $x$, average $N$ over different $x$, and define $L=L_{y}/\langle N\rangle$, where $L_{y}$ is the system size along $y$. Fig. 3(a) shows $L$ versus time by a log-log plot in the interval of $\log_{10}(L/\xi)\in(1,1.2)$, from which we estimate the exponent $p$ by a linear fit $\log_{10}\left(L/\xi\right)=p\log_{10}(|\alpha|\gamma t)+c_{0}$. The estimate gives $p=0.45$ for the 2D interface of a 3D system. The discrepancy from the theoretical result may be because the simulation has not entered the long-time region.
One can define a structure factor $S(\bm{K},t)$ on the interface by a Fourier transform of $C(\bm{R},t)$ [36, 37], which may be experimentally measurable by time-resolved X-ray diffraction [43, 37]. The correlation length $L(t)$ could be fitted from the dynamical scaling form $S(\bm{K},t)=g(KL)$  [30].
Figure 3: Numerically computed $\log_{10}\left(L/\xi\right)$ as a function of $\log_{10}(|\alpha|\gamma t)$ for $0.00012\zeta=$ 0.0001, and the linear fitting. In the numerical simulation, space and time are discretized by step sizes $0.5\xi$ and $0.03(|\alpha|\gamma)^{-1}$. Data points are selected within $\log_{10}(L/\xi)\in(1,1.2)$. (a) The 3D $U(1)$ symmetric case. (b) The 3D $U(1)$ breaking case with $\beta/|\alpha|=0.1$. Both (a) and (b) are simulated in a $L_{x}\times L_{y}\times L_{z}$ grid with $L_{z}/\xi=60$, $L_{x}/\xi=L_{y}/\xi=1280$. (c)(d) The 2D case with the $Z_{2}$ term $\beta/|\alpha|$ being $0.001$ (c) and $0.1$ (d). The phase distribution of a topological defect is sketched in the inset of (d) with the same colorbar as Fig. 2(b,c). Both (c) and (d) are simulated in a $L_{y}\times L_{z}$ grid with $L_{z}/\xi=120$, $L_{y}/\xi=38400$.
The effects of a $Z_{2}$ term— To model the case without the $U(1)$ symmetry, we add a term $2\beta\psi_{2}^{2}$ (‘$Z_{2}$ term’) with $\beta>0$ to the free energy density in Eq. (2), reducing the $U(1)$ symmetry to a $Z_{2}\otimes Z_{2}$ symmetry under the transformation $\psi_{1(2)}\rightarrow\pm\psi_{1(2)}$. Due to the $Z_{2}$ term, the low-temperature equilibrium state is long-range ordered even for a 2D system so that a well-defined 1D domain wall could be generated. This model describes commensurate CDW/SDW [44, 45] and excitonic insulators [46], etc. As shown below, the decay process of the domain wall for the small $\beta$ case is characterized by a similar two-stage dynamics as the $U(1)$ symmetric case, while the scaling behavior may be different.
For the first-stage dynamics, the $Z_{2}$ term modifies $\hat{L}_{2}$ in Eq. (3), $\hat{L}_{2}=-\xi_{0}^{2}\partial_{z}^{2}+2(\alpha+\beta)+2\psi_{0}^{2}$, so that the lowest eigenvalue of $\hat{L}_{2}$ becomes $\alpha+2\beta$. For $\beta>|\alpha|/2$, the eigenvalue is positive, and the domain wall is stable against any fluctuation of the imaginary part of $\psi$. For $\beta<|\alpha|/2$, the mode $\phi_{2}^{\bm{K}}(\bm{r})=\text{sech}(z/\xi)e^{i\bm{K}\cdot\bm{R}}$ with $K<\sqrt{|\alpha|-2\beta}/\xi_{0}$ grows exponentially in time. The growth is characterized by the correlation function $C(\bm{R},t)$. The correlation function in the long-time regime with $\gamma t\gg 1/\beta$ and $\gamma t\gg 1/(|\alpha|-2\beta)$ is computed from Eq. (3) as [29]
$C(\bm{R},t)\approx c_{1}\zeta|\alpha|\frac{e^{2\left(|\alpha|-2\beta\right)\gamma t}}{\sqrt{8\pi|\alpha|\gamma t}^{D-1}}e^{-\frac{R^{2}}{8\xi_{0}^{2}\gamma t}}$ (6)
where $c_{1}\approx\sqrt{|\alpha|/\beta}$ if $\beta\ll|\alpha|$. Similar to Eq. (4), one has a growing correlation length $L(t)\sim\xi_{0}\sqrt{8\gamma t}$. The crossover time to the second-stage dynamics is $\gamma t_{\text{c}}\sim(2|\alpha|)^{-1}\ln(\zeta^{-1})$ for $\beta\ll|\alpha|$, given that $\ln(\zeta^{-1})\gg|\alpha|/\beta$ such that $t_{\text{c}}$ is in the long-time regime.
The coarsening dynamics of the topological defects with the $Z_{2}$ term is also described by the balance equation Eq. (5), while the friction coefficient $\lambda(L)$ and the driving force $F_{\text{d}}(L)$ depends on a spatial profile of the topological defect that is deformed by the $Z_{2}$ term. In the case of $\beta\ll|\alpha|$, the free energy density $f$ is dominated by its phase part,
$f\approx|\alpha|\left(\xi_{0}^{2}|\nabla\theta|^{2}+2\beta\sin^{2}\theta\right)\,.$ (7)
Even with the small $\beta$, the topological defect is no longer a perfect circular vortex, but connects two different sine-Gordon solitons $\theta_{0}^{\pm}=\text{arg}\left(\sinh(\sqrt{2\beta}z/\xi_{0})\pm i\right)$  [29, 47], which is sketched in the inset of Fig. 3(d). When the distance $r$ from the vortex core in the 2D plane is much smaller than a length scale $\xi_{0}/\sqrt{2\beta}$, the spatial profile approaches the form of a circular symmetric vortex. For $r\gg\xi_{0}/\sqrt{2\beta}$, the profile $\theta_{0}^{\pm}$ approaches the sine-Gordon solitons as $\theta\approx\theta_{0}^{\pm}+s(\bm{r})$, where $s(\bm{r})\sim e^{-r\sqrt{2\beta}/\xi_{0}}$ decays exponentially.
As a result, the coarsening dynamics crossovers between two asymptotic time domains. In the early-time domain with $L(t)\ll\xi_{0}/\sqrt{2\beta}$, adjacent defects see each other as circular symmetric vortices. Therefore, the driving force and frictional force are the same as the $U(1)$ symmetric case. In 2D systems, they are $F_{\text{d}}(L)\sim E_{\text{c}}|\alpha|\xi_{0}^{2}/L$, and $\lambda(L)\sim E_{\text{c}}|\alpha|\ln(L/\xi)/\gamma$, respectively. From Eq. (5), one has $L(t)\sim\xi_{0}\sqrt{\gamma t/\ln(|\alpha|\gamma t)}$, which is nearly diffusive in the large-$\gamma t$ limit. In 3D systems, the calculation is explained below Eq. (5). In the late-time domain with $L(t)\gg\xi_{0}/\sqrt{2\beta}$, the adjacent vortices see each other as the sine-Gordon solitons. Thereby, the friction coefficient becomes a constant $\lambda\sim E_{\text{c}}|\alpha|/\gamma$. The driving force in the 2D system becomes short-range attraction between vortices and antivortices with $F_{\text{d}}(L)\sim\xi_{0}E_{\text{c}}|\alpha|\sqrt{2\beta}e^{-L\sqrt{2\beta}/\xi_{0}}$, while the driving force in the 3D system is dominated by the string tension with $F_{\text{d}}(L)\sim\xi_{0}^{2}E_{\text{c}}|\alpha|L^{-1}$, see Sec. III B of Ref. [29]. This results in the following scaling for the late-time domain of the coarsening dynamics,
$L(t)\sim\begin{cases}\frac{\xi_{0}}{\sqrt{2\beta}}\ln(2\beta\gamma t)&\text{in 2D systems,}\\ \xi_{0}\sqrt{\gamma t}&\text{in 3D systems.}\end{cases}$ (8)
Thus, the coarsening dynamics in the 2D systems exhibit a diffusive-to-subdiffusive crossover as $L$ goes from $L\ll\xi_{0}/\sqrt{2\beta}$ to $L\gg\xi_{0}/\sqrt{2\beta}$, while that in the 3D systems are always diffusive.
Numerical simulations of Eq. (1) with the $Z_{2}$ term shows consistent coarsening dynamics. In Fig. 3(b,c,d), we fit the value of $p$ in the interval of $\log_{10}(L/\xi)\in(1,1.2)$ for several $\beta/|\alpha|$. We get $p=0.46$ in the 3D systems with $\beta/|\alpha|=0.1$ (Fig. 3(b)), which is consistent with the diffusive dynamics. In the 2D systems, we obtain $p=0.28$ and $0.17$ with $\beta/|\alpha|=0.001$ ($L<\xi_{0}/\sqrt{2\beta}$) and $0.1$ ($L>\xi_{0}/\sqrt{2\beta}$), respectively (Fig. 3(c,d)), which is consistent with the diffusive-to-subdiffusive crossover.
Stability of the interface— In addition to the motion within the interface, the defects also move in the $z$-direction due to the thermal noise $\eta$. From Eq. (1), the $z$-direction motion of a defect is a Brownian motion given by $d(z/\xi_{0})/d(\gamma t)=H_{z}$ with a random force $H_{z}$ dependent on the defect size. For a typical defect with size $L$, the random force for a vortex satisfies
$\langle H_{z}(t)H_{z}(t^{\prime})\rangle=\begin{cases}\frac{2\zeta}{h_{z}(L)}\delta(\gamma t-\gamma t^{\prime})&\text{in 2D systems,}\\ \frac{2\zeta}{h_{z}(L)L/\xi}\delta(\gamma t-\gamma t^{\prime})&\text{in 3D systems.}\end{cases}$ (9)
Here $h_{z}(L)=\int d^{2}\bm{r}\ (\partial_{z}\theta)^{2}$ where $\theta(\bm{r})$ is the profile of a vortex with size $L$ discussed below Eq. (7), see Sec. III B of Ref. [29]. For the $U(1)$ symmetric case as well as for the $L\ll\xi_{0}/\sqrt{2\beta}$ case with a nonzero $\beta$, one has $h_{z}(L)\sim\ln(L/\xi)$, while it is hard to evaluate $h_{z}(L)$ for the other cases. Eq. (9) implies that as long as $\zeta$ is small, the $z$-direction motion of the defects is negligible and the interface is stable.
Effects of the inertia term— So far, we have clarified the decaying process of the domain wall in terms of the “model A” dynamics, Eq. (1). The “model A” dynamics neglects the inertia of the order parameter, the $\partial_{t}^{2}\psi$ term, while the inertia is crucial for the inversion of the order parameter by the pump pulse that generates the domain wall [25, 26, 27, 28, 29]. Thus, we now proceed to add the inertia term $\gamma_{0}^{-2}\partial_{t}^{2}\psi$ to the left-hand side of Eq. (1), and discuss how the decaying processes of the domain wall are modified. Note that microscopic derivations of the relevant TDGL theory [40, 48] gives $\gamma,\,\gamma_{0}\sim T$ for pure electronic density-wave orders arising from Fermi surface nesting, see Ref. [40] and Appendix E of Ref. [48], while a mixing with lattice distortions may significantly increase the inertia.
In the first-stage dynamics, a small $|\alpha|$ suppresses the inertia term more effectively than the damping term. Thus, as long as the temperature is close to $T_{\text{c}}$ where $|\alpha|\ll 1$, the inertia term does not change the qualitative picture of the first-stage dynamics, modifying only the coefficients. Specifically, the time-dependent correlation length in Eqs. (4,6) is renormalized by a factor $1/\sqrt{\kappa}$ with $\kappa=\sqrt{1+4(|\alpha|-2\beta)\gamma^{2}/\gamma_{0}^{2}}$, see Sec. IV of Ref. [29].
In the second-stage dynamics, the $\partial_{t}^{2}\psi$ term gives the topological defect an inertial mass $m(L)=\lambda(L)\gamma/\gamma_{0}^{2}$ [49], which is dependent on the defect size $L$. Thereby, Eq. (5) is modified into
$m(L)\frac{d^{2}L}{dt^{2}}+\lambda(L)\frac{dL}{dt}=F_{\text{d}}(L),$ (10)
see Sec. IV of Ref. [29]. Note that all the results of $L(t)$ satisfying Eq. (5) obtained in this work satisfy Eq. (10) in the long-time limit, when the term $m(L)d^{2}L/dt^{2}$ is much smaller than the other two terms.
Discussion—Our theory enriches the mechanism of topological defect generation in ultrafast experiments [50, 51, 52, 11, 24, 53]. It applies to a wide class of systems with order parameters breaking a (quasi) continuous symmetry, such as charge [54, 55, 45, 22, 56], spin [41], excitonic order [57, 58, 46], and superconductivity [39, 42, 40, 16].
In some experiments for 3D systems, a strong pump fluence could generate multiple domain walls at different depths in the $z$ direction with a typical inter-wall distance $L_{z}$ [26, 25, 27]. In this paper, we focused on the decaying dynamics of an individual domain wall neglecting the interaction between the walls. This is appropriate as long as the inter-wall distance $L_{z}$ is much larger than the typical size $L$ of the topological defects on the walls, such that each wall evolves independently with a diffusive growth of in-plane correlation length $L$ while $L_{z}$ does not grow. In other words, the measured correlation-length dynamics depends on the direction of the measurement. This might be relevant to a recent pump-probe experiment that reports anomalous subdiffusive coarsening dynamics of the CDW in $\text{LaTe}_{3}$ [36].
Similar physics also applies to systems with two competing orders [16, 29]. For example, a real field $\psi$ breaking a $Z_{2}$ symmetry (e.g., commensurate CDW) competes with a complex field $\Psi=\Psi_{1}+i\Psi_{2}$ breaking a $U(1)$ symmetry (e.g., superconductivity or incommensurate CDW) [55, 54] such that $\psi$ dominates and $\Psi$ is completely suppressed in the equilibrium. If a pump pulse generates a domain wall of $\psi$, the fluctuations of $\Psi$ may become unstable on the domain wall and grow into a random domain structure with increasing correlation length, see Sec. V of Ref. [29].
Acknowledgment— We thank Andrew J. Millis, Mariano Trigo, Alfred Zong, Ke Liu, Yeyang Zhang, and Zhenyu Xiao for helpful discussions. Lingxian Kong and Ryuichi Shindou acknowledge the supports from the National Basic Research Programs of China (No. 2019YFA0308401) and the National Natural Science Foundation of China (No. 11674011 and No. 12074008). Zhiyuan Sun acknowledges the support of the startup grant from the State Key Laboratory of Low-Dimensional Quantum Physics at Tsinghua University.